# Spectrum of random centrosymmetric matrices; CLT and Circular law

### v2

@@ 6.1. Proof of tightness
The sum is divided by $n$ because the matrix $T_{1}$ is scaled by $1/\sqrt{n}$. Moreover, the sum stops at $k-1$ because $\tilde{\Gamma}_{k}^{1}(z,\bar{\eta})$ is a constant for $i,j=k,k+1,\dots,\lceil n/2\rceil$. On the other hand, we have
$\frac{\partial\mathcal{R}_{z}(T_{1}^{(k)})_{ks}}{\partial t_{ij}^{1}}$ $=\mathcal{R}_{z}(T_{1}^{(k)})_{ki}\mathcal{R}_{z}(T_{1}^{(k)})_{js}\mathbf{1}_{\{j\neq k\}},$ $\frac{\partial\overline{\mathcal{R}_{z}(T_{1}^{(k)})_{ks}}}{\partial t_{ij}^{1}}$ $=0,$
@@
$\sum_{i,j=1}^{k-1}\left|\sum_{s=0}^{\lceil n/2\rceil}\left(\Lambda_{ks}\tilde{\Upsilon}_{sk}+\tilde{\Lambda}_{ks}\Upsilon_{sk}\right)\frac{\partial\omega_{\epsilon_{n},k}}{\partial a_{ij}}\mathbf{1}_{\left\{\left(\rho_{1}<\|T_{1}^{(k)}\|<\rho_{1}+\epsilon_{n}\right\}\right.}\right|^{2}$
@@
This completes the proof of the central limit theorem. Now, we demonstrate the calculation of the covariance kernel in Section 7. Before that, let us verify this result through a numerical simulation; see Figure 3.
Figure 3. Histogram of $\operatorname{L}^{\circ}(P_{d})$ of a $4000\times 4000$ random centrosymmetric matrix sampled 1000 times. The entries of the matrix are drawn from the Gaussian distribution, where $P_{d}(x)=2x^{3}+x^{4}$.
@@ 7. Calculation of variance
$\operatorname{Var}(\operatorname{Tr}(P_{d}(M)))=\sum_{k=1}^{d}2k|a_{k}|^{2}+o(1),$
where $P_{d}(x)=\sum_{k=1}^{d}a_{k}x^{k}$ is a polynomial of degree $d\geq 1$.
@@ 7.1. Analyzing expected product
We evaluate $\mathbb{E}\left[x_{1}x_{2}\cdots x_{n}\right]$, where $x_{i}s$ are random variables from Condition 3.3. Note that $\mathbb{E}[x_{i}^{2}]=0$ if $x_{i}$ satisfies Condition 3.3.
In the next calculations, we need to evaluate the terms of the forms $\mathbb{E}[\operatorname{Tr}(M^{k})]$ and $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]$. Upon expansion, the term $\operatorname{Tr}(M^{k})$ is represented as $\sum_{i_{1},i_{2},\ldots,i_{k}=1}^{n}x_{i_{1}i_{2}}$ $x_{i_{2}i_{3}}$cdots $x_{i_{k}i_{1}}$. Notice that here the random variables $x_{i_{1}i_{2}}$, $x_{i_{2}i_{3}}$, $\dots$ ,$x_{i_{k}i_{1}}$ are linked via their indices $i_{1},i_{2},\ldots,i_{k}$. The arrangement of these indices plays an important role in evaluating expectations. We formalize this using the notation of index chain as defined below.
@@
The index chain of the product of random variables $x_{i_{1}i_{2}}$ $x_{i_{2}i_{3}}$cdots $x_{i_{k}i_{1}}$ is the unique sequence $i_{1}\to i_{2}\to\cdots\to i_{k}\to i_{1}$ of indices.
@@
This will be relevant while calculating the $\mathbb{E}[\operatorname{Tr}M^{k}]$ or $\mathbb{E}[\operatorname{Tr}M^{k}\bar{M}^{l}].$ We shall use the notion of free indices in the remaining part of this section; in particular in the proofs of Lemma 7.4, Lemma 7.5, and Lemma 7.6.
Now, we proceed to the calculation of $\mathbb{E}[\operatorname{Tr}(M^{k})]$ and $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]$, which are discussed in Sections 7.2 and 7.3.
@@ 7.2. Single Chain Expectation
$\mathbb{E}[\mathrm{Tr}(M^{k})]=\frac{1}{n^{{k/2}}}\sum_{i_{1},i_{2},\dots,i_{k}=1}^{n}\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k-1}i_{k}}x_{i_{k}i_{1}}].$
@@
$\mathbb{E}[\text{Tr}(M^{k})]=\frac{1}{n^{k/2}}\sum_{i_{1},i_{2},\dots,i_{k}=1}^{n}\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k-1}i_{k}}x_{i_{k}i_{1}}].$
To evaluate $\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k-1}i_{k}}x_{i_{k}i_{1}}]$, we need to determine the number of ways in which $x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k-1}i_{k}}x_{i_{k}i_{1}}$ simplifies to a product where each $x_{ij}$ may appear with some multiplicity. However, given that $\mathbb{E}[x_{ij}]=0=\mathbb{E}[x_{ij}^{2}]$ according to Condition 3.3, we should focus on the cases where more than two random variables are equal. Since the product has $k$ random variables, essentially the the task boils down to considering the integer partitions of $k$, where group sizes in each partition is at least $3$. For example, $22=10+5+4+3$ is an integer partition of $22$ having four groups, where group sizes are $10,5,4,3$
@@
Let us analyze a generic group of size $k_{l}.$ To reduce notional complexity, here we denote the group as $x_{\alpha_{1}}x_{\alpha_{2}}\cdots x_{\alpha_{k_{l}}}$, where each $\alpha_{p}$ is of the form $(i_{\alpha_{p}},i_{\alpha_{p}+1}).$ As per the construction, $x_{\alpha_{1}}=x_{\alpha_{2}}=\cdots=x_{\alpha_{k_{l}}}.$ Therefore, using the fact that $x_{ij}$s satisfy Poincaré inequality, we have
$\mathbb{E}[x_{\alpha_{1}}x_{\alpha_{2}}\cdots x_{\alpha_{k_{l}}}]=\mathbb{E}[|x_{11}|^{k_{l}}]\leq k_{l}!\left(\frac{2}{\alpha}\right)^{k_{l}/2}.$
Hence, we conclude that if we fix a partition $\sum_{l=1}^{g}k_{l}=k$, and fix random variables in each group of the partition, then
$\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k-1}i_{k}}x_{i_{k}i_{1}}]$ $\leq\prod_{l=1}^{g}k_{l}!\left(\frac{2}{\alpha}\right)^{k_{l}/2}$
$=\left(\frac{2}{\alpha}\right)^{k/2}\prod_{l=1}^{g}k_{l}!$
$\leq\left(\frac{2}{\alpha}\right)^{k/2}k!,$
where the last inequality follows from the fact that $m!n!\leq(m+n)!.$ However, if we fix a partition of $k$, but do not fix the random variables in the partition groups, then we have some choices for the indices $\alpha_{1},\alpha_{2},\ldots,\alpha_{k_{l}}$ in the product $x_{\alpha_{1}}x_{\alpha_{2}}\cdots x_{\alpha_{k_{l}}}$, which is a group of size $k_{l}$. Now, since all the random variables in this group are equal, we only have the freedom to choose $\alpha_{1}.$ Once $\alpha_{1}$ is fixed, the remaining $\alpha_{p}$s in the same group can either be $\alpha_{1}$ or $n+1-\alpha_{1}$. Hence, for a given $\alpha_{1}$, we have at most $2^{k_{l}-1}$ many possibilities of the indices $\alpha_{2},\alpha_{3},\ldots,\alpha_{k_{l}}$ for that group.
The index $\alpha_{1}=(i_{\alpha_{1}},i_{\alpha_{1}+1})$ consists of two indices. However, since this group $x_{\alpha_{1}}x_{\alpha_{2}}\cdots x_{\alpha_{k_{l}}}$ is a part of the full chain $x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k-1}i_{k}}x_{i_{k}i_{1}}$, one of the indices among $i_{\alpha_{1}},i_{\alpha_{1}+1}$ must also appear in some other group. As a result, either $i_{\alpha_{1}}$ or $i_{\alpha_{1}+1}$ must be determined by some other group. Therefore, we have the freedom to choose only one of the indices $i_{\alpha_{1}},i_{\alpha_{1}+1}$, which can be done in at most $n$ possible ways. Hence, we have an additional factor $n2^{k_{l}-1}$ in the $l$th group of the partition. And overall, we have a factor of $n^{g}2^{\sum_{l=1}^{g}(k_{l}-1)}=n^{g}2^{k-g}.$
@@
$\mathbb{E}[\text{Tr}(M^{k})]$ $=\frac{1}{n^{k/2}}\sum_{i_{1},i_{2},\dots,i_{k}=1}^{n}\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k-1}i_{k}}x_{i_{k}i_{1}}]$
$=\frac{1}{n^{k/2}}\sum_{g=1}^{k/3}\sum_{\stackrel{{\scriptstyle k_{1}+k_{2}+\cdots+k_{g}=k}}{{k_{l}\geq 3}}}n^{g}2^{k-g}\left(\frac{2}{\alpha}\right)^{k/2}k!$
$\leq\frac{1}{n^{k/2}}\left(\frac{k}{3}\right)^{k}n^{k/3}2^{2k/3}\left(\frac{2}{\alpha}\right)^{k/2}k!$
@@
for $k>4/3\alpha$. The third last inequality follows from the fact that $n^{g}2^{k-g}\leq n^{k/3}2^{2k/3}$ for $31\leq g\leq k/$ 3 and the number of partitions of $k$ into at most $k/3$ many groups is $(k/3)^{k}.$ This completes the proof. ∎
Now, we shall proceed with the calculation of the covariance between $\operatorname{Tr}M^{k}$, $\operatorname{Tr}M^{l}.$
@@ 7.3. Double chain Expectation
$\sum_{\begin{subarray}{c}k,l=1\end{subarray}}^{d}\frac{1}{n^{(k+l)/2}}\sum_{{i_{1},i_{2},\dots i_{k},j_{1},j_{2},\dots j_{l}=1}}^{n}\mathbb{E}\left[(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\cdots\bar{x}_{j_{l}j_{1}})\right].$
Here, $k$ and $l$ are positive integers denoting the lengths of the respective chains. From the Condition 3.3, we know that $\mathbb{E}[x_{ij}]=0=\mathbb{E}[x_{ij}^{2}].$ Therefore as explained in Lemma 7.4, we need to choose the indices of the chains in such such a way that each random variable occurs with a multiplicity. However, it is quite evident that if we make more random variables equal to each other within a chain or across different chains, then we lose more degrees of freedom in the index chains. So, we aim to make fewer random variables equal, preferably $2$, across two different chains. Since in this case, we have the term $\mathbb{E}[|x_{ij}|^{2}]$, which is $1$. Thus, when we have a product of two chains such as $(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\cdots\bar{x}_{j_{l}j_{1}})$, we can pair random variables from the first chain $x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k}i_{1}}$ with the second chain $\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\cdots\bar{x}_{j_{l}j_{1}}$ in the following three different ways:
@@
Case 1 (Intra chain merging): As we have seen in the Lemma 7.4, the single chain expectation is asymptotically zero. Thus, intra chain merging gives us zero asymptotic contribution.
Case 2 (Cross chain merging): To evaluate $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]$, let us consider a simple example with $k=l=5$.
@@
We shall see in the proof of Lemma 7.5, we get maximum contribution if every random variable from the first chain gets paired with a random variable from the second chain.
One way of cross-chain merging is that if $x_{i_{t}i_{l}}$ gets paired with $\bar{x}_{j_{p}j_{q}}$, then the pairing for $x_{i_{t+1}i_{l+1}}$ will be $\bar{x}_{j_{p+1}j_{q+1}}$ for all $t$ and $l$ such that $1\leq t,l\leq 5$. In other words, the random variables from the first chain get matched with the random variables in the second chain in the same order as they appear in the first chain, see Figure 4, and we demonstrate that it gives us a nonzero asymptotic contribution.
@@
For any pair of random variables $(x_{ij},\bar{x}_{kl})$, where $x_{ij}$ and $\bar{x}_{kl}$ are from the chains $x_{i_{1}i_{2}}x_{i_{2}i_{3}}x_{i_{3}i_{4}}$ $x_{i_{4}i_{5}}x_{i_{5}i_{1}}$, and $\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\bar{x}_{j_{3}j_{4}}\bar{x}_{j_{4}j_{5}}\bar{x}_{j_{5}j_{1}}$ respectively, $\mathbb{E}[x_{ij}\bar{x}_{kl}]\neq 0$ if and only if $(i,j)=(k,l)$ or $(i,j)=(n+1-k,n+1-l)$. This leads to the following possible pairings in the above example.
$\begin{array}[]{l}\left\{\begin{array}[]{ l }{j_{1}=n+1-i_{1}}\\ {j_{2}=n+1-i_{2}}\end{array}\quad\text{ or }\quad\left\{\begin{array}[]{l}j_{1}=i_{1}\\ j_{2}=i_{2}\end{array}\right.\right.\\ \\ \left\{\begin{array}[]{ l }{j_{2}=n+1-i_{2}}\\ {j_{3}=n+1-i_{3}}\end{array}\quad\text{ or }\quad\left\{\begin{array}[]{l}j_{2}=i_{2}\\ j_{3}=i_{3}\end{array}\right.\right.\\ \\ \left\{\begin{array}[]{ l }{j_{3}=n+1-i_{3}}\\ {j_{4}=n+1-i_{4}}\end{array}\quad\text{ or }\quad\left\{\begin{array}[]{l}j_{3}=i_{3}\\ j_{4}=i_{4}\end{array}\right.\right.\\ \\ \left\{\begin{array}[]{ l }{j_{4}=n+1-i_{4}}\\ {j_{5}=n+1-i_{5}}\end{array}\quad\text{ or }\quad\left\{\begin{array}[]{l}j_{4}=i_{4}\\ j_{5}=i_{5}\end{array}\right.\right.\\ \\ \left\{\begin{array}[]{ l }{j_{5}=n+1-i_{5}}\\ {j_{1}=n+1-i_{1}}\end{array}\quad\text{ or }\quad\left\{\begin{array}[]{l}j_{5}=i_{5}\\ j_{1}=i_{1}\end{array}\right.\right.\\ \end{array}$
If we start with $j_{1}=n+1-i_{1}$, then the chain has to follow the constraints listed in the left column; not the ones which are listed on the right column. Similarly, if we start with the condition $j_{1}=i_{1}$, then we have to follow the constraints that are listed on the right column.
@@
Thus, we conclude that when we do the cross chain merging for the first and the second chain, namely $i_{1}\rightarrow i_{2}\rightarrow i_{3}\rightarrow i_{4}\rightarrow i_{5}\to i_{1}$ and $j_{1}\rightarrow j_{2}\rightarrow j_{3}\rightarrow j_{4}\rightarrow j_{5}\to j_{1}$, one way for cross chain merging is to copy the first chain in place of the second chain, resulting in: $i_{1}\rightarrow i_{2}\rightarrow i_{3}\rightarrow i_{4}\rightarrow i_{5}\to i_{1}\rightarrow i_{2}\rightarrow i_{3}\rightarrow i_{4}\rightarrow i_{5}\to i_{1}$. This represents one possibility with $5$ free indices. Additionally, we can start the second chain with $i_{2}$ or $i_{3}$ or $\dots$ or $i_{5}$ like $i_{2}\rightarrow i_{3}\rightarrow\dots\rightarrow i_{5}\rightarrow i_{1}\rightarrow i_{2}$, resulting in $4$ more choices. Therefore, we have $5n^{5}$ combinations for cross chain merging of the first and the second chain so far. Moreover, since $x_{(i,j)}=x_{(n+1-i,n+1-j)}$ , we can perform a cross chain merging by changing the second chain to $n+1-i_{1}\rightarrow n+1-i_{2}\rightarrow\dots\rightarrow n+1-i_{5}\rightarrow n+1-i_{1}$, which will contribute another $5n^{5}.$ Hence, the total number of combinations for cross chain merging of the first and second chain is $2\times 5n^{5}$ from type of merging in Figure 4.
Now, let us consider the another type of case as demonstrated in Figure 5, where the first chain is not merging with the second chain in the same sequence. We shall argue that such mergings will give negligible contributions.

### v3

@@ 6.1. Proof of tightness
The sum is divided by $n$ because the matrix $T_{1}$ is scaled by $1/\sqrt{n}$. Moreover, the sum stops at k1k-1 because $\tilde{\Gamma}_{k}^{1}(z,\bar{\eta})$ is a constant for $i,j=k,k+1,\dots,\lceil n/2\rceil$. On the other hand, we have
$\frac{\partial\mathcal{R}_{z}(T_{1}^{k})_{ks}}{\partial t_{ij}^{1}}$ $=\mathcal{R}_{z}(T_{1}^{k})_{ki}\mathcal{R}_{z}(T_{1}^{k})_{js}\mathbf{1}_{\{j\neq k\}},$ $\frac{\partial\overline{\mathcal{R}_{z}(T_{1}^{k})_{ks}}}{\partial t_{ij}^{1}}$ $=0,$
@@
$\sum_{i,j=1}^{k-1}\left|\sum_{s=0}^{\lceil n/2\rceil}\left(\Lambda_{ks}\tilde{\Upsilon}_{sk}+\tilde{\Lambda}_{ks}\Upsilon_{sk}\right)\frac{\partial\omega_{\epsilon_{n},k}}{\partial a_{ij}}\mathbf{1}_{\left\{\left(\rho_{1}<\|T_{1}^{k}\|<\rho_{1}+\epsilon_{n}\right\}\right.}\right|^{2}$
@@
This completes the proof of the central limit theorem. We now demonstrate the calculation of the covariance kernel in Section 7. Before that let us verify this result through numerical simulation; see Figure 3.
Figure 3. Histogram of $\operatorname{L}^{\circ}(P_{d})/\sqrt{n}$ of a $4000\times 4000$ random centrosymmetric matrix sampled 1000 times. The entries of the matrix are drawn from the Gaussian distribution, where $P_{d}(x)=2x^{3}+x^{4}$.
@@ 7. Calculation of variance
$\operatorname{Var}(\operatorname{Tr}(P_{d}(M)))=\sum_{k=1}^{d}2ka_{k}^{2}+o(1),$
where $P_{d}(x)=\sum_{k=1}^{d}a_{k}x^{k}$ is a polynomial of degree $d\geq 1$ with real coefficients.
@@ 7.1. Analyzing expected product
We evaluate $\mathbb{E}\left[x_{1}x_{2}\ldots x_{n}\right]$, where $x_{i}s$ are random variables from condition (3.3). Note that $\mathbb{E}[x_{i}^{2}]=0$ if $x_{i}$ satisfies condition (3.3).
In the next calculations, we need to evaluate the terms of the forms $\mathbb{E}[\operatorname{Tr}(M^{k})]$ and $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]$. Upon expansion, the term $\operatorname{Tr}(M^{k})$ is represented as $\sum_{i_{1},i_{2},\ldots,i_{k}=1}^{n}x_{i_{1}i_{2}}$ $x_{i_{2}i_{3}}$ $\cdots$ $x_{i_{k}i_{1}}$. Notice that here the random variables $x_{i_{1}i_{2}}$, $x_{i_{2}i_{3}}$, $\dots$ $x_{i_{k}i_{1}}$ are linked via their indices $i_{1},i_{2},\ldots,i_{k}$. The arrangement of these indices plays an important role in evaluating expectations. We formalize this using the notation of index chain as defined below.
@@
The index chain of the product of random variables $x_{i_{1}i_{2}}$ $x_{i_{2}i_{3}}$ $\cdots$ $x_{i_{k}i_{1}}$ is the unique sequence $i_{1}\to i_{2}\to\cdots\to i_{k}\to i_{1}$ of indices.
@@
This will be relevant while calculating the $\mathbb{E}[\operatorname{Tr}M^{k}]$ or $\mathbb{E}[\operatorname{Tr}M^{k}\bar{M}^{l}].$ We shall use the notion of free indices in the remaining part of this section; in particular in the proof of Lemma 7.4, Lemma 7.5, and Lemma 7.6.
Now, we proceed by calculating $\mathbb{E}[\operatorname{Tr}(M^{k})]$ and $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]$ in two different scenarios such as single chain and double chain merging.
@@ 7.2. Single Chain Expectation
$\mathbb{E}[\mathrm{Tr}(M^{k})]=\frac{1}{n^{{k/2}}}\sum_{i_{1},i_{2},\dots,i_{k}=1}^{n}\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{(k-1)}i_{k}}x_{i_{k}i_{1}}].$
@@
$\mathbb{E}[\text{Tr}(M^{k})]=\frac{1}{n^{k/2}}\sum_{i_{1},i_{2},\dots,i_{k}=1}^{n}\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{(k-1)}i_{k}}x_{i_{k}i_{1}}].$
To evaluate $\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{(k-1)}i_{k}}x_{i_{k}i_{1}}]$, we need to determine the number of ways in which $x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{(k-1)}i_{k}}x_{i_{k}i_{1}}$ simplifies to a product where each $x_{ij}$ may appear with some multiplicity. However, given that $\mathbb{E}[x_{ij}]=0=\mathbb{E}[x_{ij}^{2}]$ according to Condition 3.3, we should focus on the cases where more than two random variables are equal. Since the product has $k$ random variables, essentially the the task boils down to considering the integer partitions of $k$, where group sizes in each partition is at least $3$. For example, $22=10+5+4+3$ is an integer partition of $22$ having four groups, where group sizes are $10,5,4,3.$
@@
Let us analyze a generic group of size $k_{l}.$ To reduce notional complexity, here we denote the group as $x_{\alpha_{1}}x_{\alpha_{2}}\ldots x_{\alpha_{k_{l}}}$, where each $\alpha_{p}$ is of the form $(i_{\alpha_{p}},i_{\alpha_{p}+1}).$ As per the construction, $x_{\alpha_{1}}=x_{\alpha_{2}}=\cdots=x_{\alpha_{k_{l}}}.$ Therefore using the fact that $x_{ij}$s satisfy Poincaré inequality, we have
$\mathbb{E}[x_{\alpha_{1}}x_{\alpha_{2}}\ldots x_{\alpha_{k_{l}}}]=\mathbb{E}[|x_{11}|^{k_{l}}]\leq k_{l}!\left(\frac{2}{\alpha}\right)^{k_{l}/2}.$
Hence we conclude that if we fix a partition $\sum_{l=1}^{g}k_{l}=k$, and fix random variables in each group of the partition, then
$\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{(k-1)}i_{k}}x_{i_{k}i_{1}}]$ $\leq\prod_{l=1}^{g}k_{l}!\left(\frac{2}{\alpha}\right)^{k_{l}/2}$
$=\left(\frac{2}{\alpha}\right)^{k}\prod_{l=1}^{g}k_{l}!$
$\leq\left(\frac{2}{\alpha}\right)^{k}k!,$
where the last inequality follows from the fact that $m!n!\leq(m+n)!.$ However, if we fix a partition of $k$, but do not fix the random variables in the partition groups, then we have some choices for the indices $\alpha_{1},\alpha_{2},\ldots,\alpha_{k_{l}}$ in the product $x_{\alpha_{1}}x_{\alpha_{2}}\ldots x_{\alpha_{k_{l}}}$, which is a group of size $k_{l}$. Now since all the random variables in this group are equal, we only have the freedom to choose $\alpha_{1}.$ Once $\alpha_{1}$ is fixed, the remaining $\alpha_{p}$s in the same group can either be $\alpha_{1}$ or $n+1-\alpha_{1}$. Hence for a given $\alpha_{1}$, we have $2^{k_{l}-1}$ many possibilities of the indices $\alpha_{2},\alpha_{3},\ldots,\alpha_{k_{l}}$ for that group.
Now the index $\alpha_{1}=(i_{\alpha_{1}},i_{\alpha_{1}+1})$ consists of two indices. However, since this group $x_{\alpha_{1}}x_{\alpha_{2}}\ldots x_{\alpha_{k_{l}}}$ is a part of the full chain $x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{(k-1)}i_{k}}x_{i_{k}i_{1}}$, one of the indices among $i_{\alpha_{1}},i_{\alpha_{1}+1}$ must also appear in some other group. As a result, either $i_{\alpha_{1}}$ or $i_{\alpha_{1}+1}$ must be determined by some other group. Therefore we have the freedom to choose only one of the indices $i_{\alpha_{1}},i_{\alpha_{1}+1}$, which can be done in at most $n$ possible ways. Hence we have an additional factor $n2^{k_{l}-1}$ in the $l$th group of the partition. And overall, we have a factor of $n^{g}2^{\sum_{l=1}^{g}(k_{l}-1)}=n^{g}2^{k-g}.$
@@
$\mathbb{E}[\text{Tr}(M^{k})]$ $=\frac{1}{n^{k/2}}\sum_{i_{1},i_{2},\dots,i_{k}=1}^{n}\mathbb{E}[x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{(k-1)}i_{k}}x_{i_{k}i_{1}}]$
$=\frac{1}{n^{k/2}}\sum_{g=1}^{k/3}\sum_{\stackrel{{\scriptstyle k_{1}+k_{2}+\cdots+k_{g}=k}}{{k_{l}\geq 3}}}n^{g}2^{k-g}\left(\frac{2}{\alpha}\right)^{k}k!$
$\leq\frac{1}{n^{k/2}}\left(\frac{k}{3}\right)^{k}n^{k/3}2^{2k/3}\left(\frac{2}{\alpha}\right)^{k}k!$
@@
for $k>4/3\alpha$. The third last inequality follows from the fact that $n^{g}2^{k-g}\leq n^{k/3}2^{2k/3}$ for $1\leq g\leq k/3$ and the number of partitions of $k$ into at most $k/3$ many groups is $(k/3)^{k}.$ This completes the proof. ∎
We shall now proceed with the calculation of the covariance between $\operatorname{Tr}M^{k}$, $\operatorname{Tr}M^{l}.$
@@ 7.3. Double chain Expectation
$\sum_{\begin{subarray}{c}k,l=1\end{subarray}}^{d}\frac{1}{n^{(k+l)/2}}\sum_{{i_{1},i_{2},\dots i_{k},j_{1},j_{2},\dots j_{l}=1}}^{n}\mathbb{E}\left[(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\dots\bar{x}_{j_{l}j_{1}})\right].$
Here, $k$ and $l$ are positive integers denoting the lengths of the respective chains. From the condition 3.3, we know that $\mathbb{E}[x_{ij}]=0=\mathbb{E}[x_{ij}^{2}].$ Therefore as explained in Lemma 7.4, we need to choose the indices of the chains in such such a way that each random variable occurs with a multiplicity. However, it is quite evident that if we make more random variables equal to each other within a chain or across different chains, then we lose more degrees of freedom in the index chains. Therefore, we aim to make fewer random variables equal, preferably $2$, across two different chains. Since in this case, we have the term $\mathbb{E}[|x_{ij}|^{2}]$, which is $1$. Thus, when we have a product of two chains such as $(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\dots\bar{x}_{j_{l}j_{1}})$, we can pair random variables from the first chain $x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{k}i_{1}}$ with the second chain $\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\dots\bar{x}_{j_{l}j_{1}}$ in the following three different ways:
@@
Case 1 (Intra chain merging): As we have seen in the Lemma 7.4, the single chain expectation is asymptotically zero. Therefore, intra chain merging gives us zero asymptotic contribution.
Case 2 (Cross chain merging): Let us consider a simple example to evaluate $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]$ with $k=l=5$. Thus,
@@
As we shall see in the proof of Lemma 7.5, we get maximum contribution if every random variable from the first chain gets paired with a random variable from the second chain.
One way of cross chain merging is illustrated in Figure 4, and we demonstrate that it gives us a nonzero asymptotic contribution.
@@
For any pair of random variables $(x_{ij},\bar{x}_{kl})$, where $x_{ij}$ and $\bar{x}_{kl}$ are from a chain $x_{i_{1}i_{2}}x_{i_{2}i_{3}}x_{i_{3}i_{4}}$ $x_{i_{4}i_{5}}x_{i_{5}i_{1}}$, and $\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\bar{x}_{j_{3}j_{4}}\bar{x}_{j_{4}j_{5}}\bar{x}_{j_{5}j_{1}}$ respectively, $\mathbb{E}[x_{ij}\bar{x}_{kl}]\neq 0$ if and only if $(i,j)=(k,l)$ or (i,j)=(n+1k,n+1l)(i,j)=(n+1-k,n+1-l). This leads to the following possible pairings in the above example.
[…]
If we start with $j_{1}=n+1-i_{1}$, then the chain has to follow the constraints listed in the left column; not the ones which are listed on the right column. Similarly, if we start with the condition $j_{1}=i_{1}$, then we have to follow the constraints that are listed on the right column. Mixing the conditions listed in the left and right column will reduce the number of free indices, leading to an asymptotically zero contribution, as the full term is divided by $n^{5}$.
@@
Thus, we conclude that when we do the cross chain merging for the first and the second chain, namely $i_{1}\rightarrow i_{2}\rightarrow i_{3}\rightarrow i_{4}\rightarrow i_{5}\to i_{1}$ and $j_{1}\rightarrow j_{2}\rightarrow j_{3}\rightarrow j_{4}\rightarrow j_{5}\to j_{1}$, one way for cross chain merging is to copy the first chain in place of the second chain, resulting in: $i_{1}\rightarrow i_{2}\rightarrow i_{3}\rightarrow i_{4}\rightarrow i_{5}\to i_{1}\rightarrow i_{2}\rightarrow i_{3}\rightarrow i_{4}\rightarrow i_{5}\to i_{1}$. This represents one possibility with $5$ free indices. Additionally, we can start the second chain with $i_{2}$ or $i_{3}$ or $\dots$ or $i_{5}$ like $i_{2}\rightarrow i_{3}\rightarrow\dots\rightarrow i_{5}\rightarrow i_{1}\rightarrow i_{2}$, resulting in $4$ more choices. Therefore, we have $5n^{5}$ combinations for cross chain merging of the first and the second chain so far. Moreover, since $x_{(i,j)}=x_{(n+1-i,n+1-j)}$ , we can perform a cross chain merging by changing the second chain to $(n+1-i_{1})\rightarrow(n+1-i_{2})\rightarrow\dots\rightarrow(n+1-i_{5})\rightarrow(n+1-i_{1})$, which will contribute another $5n^{5}.$ Therefore, the total number of combinations for cross chain merging of the first and second chain is $2\times 5n^{5}$ from type of merging in Figure 4.
So far, we have observed that if $x_{i_{t}i_{l}}$ finds its pair $\bar{x}_{j_{p}j_{q}}$, then the pairing for $x_{i_{t+1}i_{l+1}}$ is $\bar{x}_{j_{p+1}j_{q+1}}$ for all $t$ and $l$ such that $1\leq t,l\leq 5$. In other words, the random variables from the first chain match with the random variables in the second chain in the same order as they appear in the first chain. Now, let us consider the another type of case as demonstrated in Figure 5, where the first chain is not merging with the second chain in the same sequence. We shall argue that such mergings will give negligible contributions.